A dead body was found within a closed room of a house where the temperature was a constant 22° C. At the time of discovery the core temperature of the body was determined to be 27° C. One hour later a second measurement showed that the core temperature of the body was 25° C. Assume that the time of death corresponds to t = 0 and that the core temperature at that time was 37° C. Determine how many hours elapsed before the body was found. [Hint: Let  t1 > 0 denote the time that the body was discovered.] (Round your answer to one decimal place.)

____ hr
Answer #1

front of di/dt tell that flee thimper Gian thou MO con ese necotons low of cooling to determine when the body was givern thaIn 192-1) =-ck+tc) wring properties of logg we can write elet lette 22-T LA - -ce 2-T the reciprocal of both sides How taking27 2) IS at feu fivewean 5 = 2 +15 pill(t+1) looking sobivalt 22 Korom both sides - 15 ekt okt KU log face tou natural of bo-1.0986122 -k : kohted + RR 5-1.0986122 +1.60043 c=0-5108178 no tow we have k our eauation for the teuperabre retary fint i